OpenStudy (anonymous):

if g^-1(x) is the inverse of g(x) what is the value of g^-1(g(x))

OpenStudy (bibby):

let's use an example g(x)=x-2 y=x-2 x=y-2 y=x+2 so g(x) = x-2 \(g^{-1}(x)\)=x+2 \(g^{-1}(g(x))=(x-2)+2=x\)

OpenStudy (bibby):

basically, whenever you pass the original to the inverse, you'll get x. that's the point of it being the inverse. it's like addition/subtraction and multiplication/division

OpenStudy (anonymous):

so it would be x then?

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

yeah, probably, maybe

OpenStudy (bibby):

for reference, \(g^{-1}(g(x))=g(g^{-1}(x))=x\)

OpenStudy (anonymous):

thanks @bibby

OpenStudy (bibby):

any time :D